Using the Nernst equation, determine the concetration of Cu(II) needed to produce a concentration cell, in combination with the provided 0.10 M Cu(II) solution, that will have a measured cell potential of 0.030 V.
Answer #1

Cu2+ +2e- > Cu

Eo for above reaction is = 0.34 V

E cell = Eo cell - (0.059/n) log (c1/c2)

0.03 = 0.34 - (0.059/2) log (c1/c2)

log (c1/c2) = 10.51

c1/c2 = 3.22*10^10

There are 2 possible possible answers depending on the avialable call of 0.1 M is used as cathode or anode.

If c1= 0.1 M then c2= 3.11*10^-12 M

If c1= 0.1 M then c2= 3.22*10^9 M

Answers are:

3.11*10^-12 M or 3.22*10^9 M